# Is a Dissipative System Always a Gradient System or a Gradient Like System?

### v1

@@ Abstract
We find that to the dynamics of a given dissipative system a p=1 differential form can be associated with a general decomposition into a potential term and a non-potential residual part. If the residual part is absent the form is closed and the system is gradient system or gradient like. If it is non-closed, in the differential form approach, it remains non-closed under a variable change of coordinates, i.e., the system is not a gradient one or a gradient like in any coordinate system. On the other hand, there are claims that a potential should always exists, i.e., the class of dissipative systems and the the class of gradient systems should coincide. We fix this conundrum by introducing a generalized change of coordinates that aims a transformation to a gradient system or a gradient like system. The condition of being closed in the new coordinates of a certain, through the generalized change of coordinates defined differential form, results in a nonlinear differential equation together with a consistency condition. We give examples of physical systems where an analytical solution for the transformation can be found, and hitherto, the potential, but even when the potential is not accessible analytically, we find that it always exist, and therefore we give in principle an affirmative answer to the defining question of this work. Our findings removes loopholes in the question if a potential may exist but it is not known.
@@
Differential forms [13][14][15] have being used in expressing in a natural way fundamental theories like the Maxwell equations for electromagnetism or the Einstein general theory of relativity [16][17] (see [19]). Differential forms are basic for defining the symplectic structure of classical Hamiltonian mechanics[20]. For our purposes we introduce the essentials needed in the sequel, and adhere to the conviction in [18].
@@ IV.1 Two-dimensional Josephson Junctions Arrays:
In this work, two relevant results are obtained. On the one hand, from the theory of differential forms, we show that a p=1 differential form can be associated with a dissipative systems. This diff-form may be closed or not. If it is a closed form then we can associate to the dissipative system a potential. it constitutes a gradient system or a gradient like system. If it is no close, (Eqn.(14), there exist no coordinate transformation that could change this property. The property of been non-closed is invariant under coordinate transformation. Through the introduction of the Homotopic or k-operator, in combination with the d-operator, in either case, one can decompose the diff-form in an exact pat and an antiexact part. This decomposition is universal valid, and express that to any dissipative systems, its dynamics is given by a potential like part and a non-potential term. On the other hand, If it is not a closed form, we proposed in this work , a nonlinear matching procedure by ”rotating” both, the vector function $\mathcal{G}(\vec{y})$ and the coordinate system $\vec{y}$, simultaneously under the transformation $\phi(\vec{y})$ and imposing for the new p=1 diff-form the necessary condition for being closed. This is the generalized change of variables (Eqn.(22) which we implement through the introduction of a matrix function (Eqn.(23). For the p=1 diff-form constructed (Eqn.(23), we derive an equation (Eqn.(26)) whose results give the equation defining this transformation, together with the consistency condition (Eqn.(27). It constitutes our main result. We claim that the same conditions that guaranties the existence of a solution of the dynamical systems (Eqn.(9)), also guaranties the solution a Eqn.(26). For some cases, the transformation can be defined through a constant matrix, and an analytical solution can be found, i.e., the case of two dimensional Josephson arrays under a magnetic field [35], [36]. For the cases in [43], and [34], it appears to us that only a numerical study is feasible. These results bring us to the main purpose of this work, namely, to answer the question, is a dissipative systems always a gradient system ?. The examples discussed above, show certainly, that whereas the differential form formalism say there is not a potential, our approach finds that indeed it exist, in the context of the generalized change of variable concept. For the time being, we can only say for sure, that for the kind of systems where Eqt.(30) apply, one can find a potential. Also, for systems where Eqt.(33) applies, although the resulting algebraic equations appears to be accessible only through pure numerical simulations. And in general, depending on each particular system, one has to figure out a strategy for find the generalized transformation defined in Then we have the natural question: given a particular dissipative systems there exist such a transformation $\phi(\vec{y})$, as defined in Eqn.(23), , and such that it satisfies Eqn.(25), Eqn.(26), and Eqn.(27). This procedure may be a difficult one, but even in principle, all dissipative systems are gradient systems , or gradient like systems. For extended systems represented by an evolution equation like a parabolic partial differential equations, the corresponding Banach space is of infinite dimension, and the associated operators are better treated in a Hilbert space with a combination of the concepts of differential forms and functional analysis. This matter is under investigation. Summarizing, our approach to the problem to the existing of a potential for a dissipative dynamical systems provides a clear mathematical set up for finding the potential, and confirms the equality of the class of dissipative systems and the class of gradient systems.
@@ VI Apendix
We thank Prof. Dr. Thomas Klinker, Universitaet Hamburg, Germany, for donating the important book for this work on Differential Forms by Henry Cartan and for financial support for this project. We thank Dr. Hector Giusti, DIC USB, for donating a PC which was used exclusively for writing this work.

### v2

@@ Abstract
We find that to a given dissipative system a p=1 differential form can be associated which has a general decomposition into a potential term and a non-potential residual part. If the residual part is absent the form is closed and the system is gradient system or gradient like. If it is non-closed, in the differential form approach, it remains non-closed under a variable change of coordinates, i.e., the system is not a gradient one or a gradient like in any coordinate system. On the other hand, there are claims that a potential should always exists, i.e., the class of dissipative systems and the the class of gradient systems should coincide. We fix this conundrum by introducing a generalized change of coordinates that aims a transformation to a gradient system or a gradient like system. The condition of being closed in the new coordinates of a certain, through the generalized change of coordinates defined differential form, results in a nonlinear differential equation together with a consistency condition. We give examples of physical systems where an analytical solution for the transformation can be found, and hitherto, the potential, but even when the potential is not accessible analytically, we find that it always exist, and therefore we give in principle an affirmative answer to the defining question of this work. Our findings removes loopholes in the question if a potential may exist but it is not known.
@@
Differential forms [13][14][15] have being used in expressing in a natural way fundamental theories like the Maxwell equations for electromagnetism or the Einstein general theory of relativity [16][17] (see [19]). Differential forms are basic for defining the symplectic structure of classical Hamiltonian mechanics[20]. For our purposes we introduce the essentials needed in the sequel, and adhere to the conviction in [18] that this rather short introduction will position the reader in catching the main points of our approach.
@@ IV.1 Two-dimensional Josephson Junctions Arrays:
In this work, two relevant results are obtained. On the one hand, from the theory of differential forms, we show that a p=1 differential form can be associated with a dissipative systems. This diff-form may be closed or not. If it is a closed form then we can associate to the dissipative system a potential. it constitutes a gradient system or a gradient like system. If it is no close, (Eqn.(14), there exist no coordinate transformation that could change this property. The property of been non-closed is invariant under coordinate transformation. Through the introduction of the Homotopic or k-operator, in combination with the d-operator, in either case, one can decompose the diff-form in an exact pat and an antiexact part. This decomposition is universal valid, and express that to any dissipative systems, its dynamics is given by a potential like part and a non-potential term. On the other hand, If it is not a closed form, we proposed in this work , a nonlinear matching procedure by ”rotating” both, the vector function $\mathcal{G}(\vec{y})$ and the coordinate system $\vec{y}$, simultaneously under the transformation $\phi(\vec{y})$ and imposing for the new p=1 diff-form the necessary condition for being closed. This is the generalized change of variables (Eqn.(22) which we implement through the introduction of a matrix function (Eqn.(23). For the p=1 diff-form constructed (Eqn.(23), we derive an equation (Eqn.(26)) whose results give the equation defining this transformation, together with the consistency condition (Eqn.(27). It constitutes our main result. We claim that the same conditions that guaranties the existence of a solution of the dynamical systems (Eqn.(9)), also guaranties the solution a Eqn.(26). For some cases, the transformation can be defined through a constant matrix, and an analytical solution can be found, i.e., the case of two dimensional Josephson arrays under a magnetic field [35], [36]. For the cases in [43], and [34], it appears to us that only a numerical study is feasible. These results bring us to the main purpose of this work, namely, to answer the question, is a dissipative systems always a gradient system ?. The examples discussed above, show certainly, that whereas the differential form formalism say there is not a potential, our approach finds that indeed it exist, in the context of the generalized change of variable concept. For the time being, we can only say for sure, that for the kind of systems where Eqt.(30) apply, one can find a potential, i.e., the two-dimensional Josephson arrays. Also, for systems where Eqt.(33) applies, i.e., Josephson junction circuits, although the resulting algebraic equations appears to be accessible only through pure numerical simulations. And in general, depending on each particular system, one has to figure out a strategy for find the generalized transformation defined in Then we have the natural question: given a particular dissipative systems there exist such a transformation $\phi(\vec{y})$, as defined in Eqn.(23), , and such that it satisfies Eqn.(25), Eqn.(26), and Eqn.(27). This procedure may be a difficult one, but even in principle, all dissipative systems are gradient systems , or gradient like systems. Our work naturally conduct us to consider dissipative systems defined in infinite dimensional Bach spaces, for example, Navier-Stokes equations [45], or the Generalized Time-dependent Ginzburg-Landau equations (GTDGLE)[46],[47],[48]. This is an extended systems represented by an evolution equation given by an abstract parabolic partial differential equations. The corresponding Banach space is of infinite dimension, and the associated operators are treated in a Hilbert space with a combination of the concepts of differential forms and functional analysis. In analogy with finite dimensional dissipative systems[49], for example the GTDGLE posses attractors with fractal dimension in Banach Spaces[50]. This matter is under investigation. Summarizing, our approach to the problem to the existing of a potential for a dissipative dynamical systems provides a clear mathematical set up for finding the potential, and confirms the equality of the class of dissipative systems and the class of gradient systems.
@@ VI Apendix
We thank Prof. Dr. Thomas Klinker, Universitaet Hamburg, Germany, for donating the important book for this work on Differential Forms by Henry Cartan and for financial support for this project. I thank Dr. Hector Giusti , DIC.USB, for donating a pc, which was used for writing this work.